APPLY THE CONCEPTS: Calculate the break-even point in sales dollars for Lennon Products Further analysis of Lennon Products’s fixed costs revealed that the company actually faces annual fixed overhead costs of $9,800 and annual fixed selling and administrative costs of $4,200. Variable cost estimates are correct: direct materials cost, $4.00 per unit; direct labor costs, $5.00 per unit; and variable overhead costs, $1.00 per unit. At this time, the selling price of $20 will not change. Complete the following formulas for the revised fixed costs. Enter the ratio as a percentage. Contribution Margin per Unit = $ – $ = $ Contribution Margin Ratio = $ = % $ Now complete the formulas for (1) the break-even point in sales dollars and (2) the units sold at the break-even point. To calculate this, divide the break-even point in sales dollars by the unit selling price. Break-Even Point in Sales Dollars = $ = $ % Units Sold at Break-Even Point = units Assume that the number of units that Lennon sold exceeded the break-even point by one (1). How much would operating income be? $ What would operating income be if the units sold exceeded the break-even point by five (5) units? $
Answer #1

Breakeven point in sales dollars=total fixed costs/contribution margin ratio
=(9800+4200)/50%
=28000
Units sold=fixed costs/contribution margin
=(9800+4200)/10=1400
Operating income:
=sales-variable costs-fixed costs
=(20*1401)-(10*1401)-(9800+4200)
=10
5 units:
=(20*1405)-(10*1405)-(9800+4200)
=50